how many moles of h2o contain 1kg of electron?

Grade:11

1 Answers

Pankaj
8 years ago
Number of electrons in one mole H2O = 10 mole
= 0 x 6.023 x 1023 electrons = 6.023 x 1024 electrons

Mass of one electron 9.1 x 10-31 kg
or
Mass of6.023 x 1024 electrons =6.023 x 1024 x 9.1 x 10-31 kg = 54.8 x 10-7 kg
or
54.8 x 10-7 kg = 1 mole H2O
1 kg = 1/54.8 x 10-7= 182481 molecules